jude's fish tank holds 100000cm^3 of water when full jude decides to pour in water to a height of 5 cm below the top of the tank how much water in cm^3 will jude need to pour into tank so that the water is 5 cm below the top?
width 40 cm depth 1 meter

Answer by mananth(16949)

jude's fish tank holds 100000cm^3 of water when full jude decides to pour in water to a height of 5 cm below the top of the tank how much water in cm^3 will jude need to pour into tank so that the water is 5 cm below the top?
width 40 cm depth 1 meter
1m = 100 cm
...
Volume of Tank = L*W*H
H= 100000/40*100
H= 25 cm the height
..
5 cm less than the height
=20 cm
..
V20 = 40*100*20=80000 cm^3
...
m.ananth@hotmail.ca